OpenStudy (anonymous):

Find the inverse function f(x)=x+9

OpenStudy (ksaimouli):

y=x+9

OpenStudy (ksaimouli):

switch x and y

OpenStudy (ksaimouli):

x=y+9

OpenStudy (ksaimouli):

and now slove for y

OpenStudy (ksaimouli):

what did u get?

OpenStudy (anonymous):

-9?

OpenStudy (ksaimouli):

solve for y =x-9

OpenStudy (ksaimouli):

the answer is x-9

OpenStudy (anonymous):

oh so you keep the x wait what if it was like f(x)=3x+6

OpenStudy (ksaimouli):

then switch again y=3x+6 ->x=3y+6->(x-6)/3=y

OpenStudy (anonymous):

how did u get that?

OpenStudy (ksaimouli):

just switch x and y and slove for y

OpenStudy (ksaimouli):

as earlier

OpenStudy (anonymous):

thanks so much :)

OpenStudy (ksaimouli):

:-)